OpenStudy (anonymous):

What is 2 to the power of 3 over 2 equal to?

OpenStudy (alexandervonhumboldt2):

2^3/2?

OpenStudy (alexandervonhumboldt2):

^2sqrt(2^3)

OpenStudy (anonymous):

Yes @AlexandervonHumboldt2

OpenStudy (anonymous):

4

OpenStudy (alexandervonhumboldt2):

no direct answers

OpenStudy (anonymous):

@yyy Please use the Code Of Conduct rules of Openstudy. Dont give direct answers.

OpenStudy (alexandervonhumboldt2):

and how did yu got that?

OpenStudy (anonymous):

@AlexandervonHumboldt2 4 Is not even one of the Answer choices.

OpenStudy (alexandervonhumboldt2):

sqrt(2^3)=2sqrt2

OpenStudy (alexandervonhumboldt2):

i didnt tell 4 yyy told 4

OpenStudy (anonymous):

I think the answer is ^3 To the square root of 16? Is this right by any chance?

OpenStudy (alexandervonhumboldt2):

2sqrt2

OpenStudy (alexandervonhumboldt2):

because \[\sqrt[2]{2^3}\]

OpenStudy (anonymous):

See, to get a no. out of the root, you need the no. that many times inside the root. Here, 2^3 has three twos. and you have a square root attached. Hence 1 two can come out, while 1 two still remains in the root. so the anser might be 2 * sq. root of 2

OpenStudy (alexandervonhumboldt2):

\[=\sqrt8=2\sqrt2\]

OpenStudy (anonymous):

Yeah if @AlexandervonHumboldt2's is the expression, then I think the answers' right :-D :-) *And the approach in that case :P

OpenStudy (anonymous):

Answer choices are: square root of 8 cube root of 8 cube root of 16 square root of 16

OpenStudy (anonymous):

square root of 8 it is :-)

OpenStudy (alexandervonhumboldt2):

teah it is just not a simplified form sqrt8

OpenStudy (anonymous):

So it would be ^2 sqrt of 8?

OpenStudy (alexandervonhumboldt2):

yes but is writtes as sqrt8

OpenStudy (anonymous):

Thanks guys i really appreciate that @AlexandervonHumboldt2 @Abhilash11

OpenStudy (anonymous):

May you help me with 1 more maybe?

OpenStudy (anonymous):

Sure man. Not a problem :-)

OpenStudy (anonymous):

Yeah, I'd try :-D

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − 2y = 2 3x + y = 6

OpenStudy (anonymous):

x-2y=2 3x+y=6 What is the value of x in the solution to the following system of equations?

OpenStudy (alexandervonhumboldt2):

elimination is the best way

OpenStudy (alexandervonhumboldt2):

x-2y+3x+y=8 -y+4x=8

OpenStudy (alexandervonhumboldt2):

y=4x-8 y=6-3x

OpenStudy (alexandervonhumboldt2):

6-3x=4x-8

OpenStudy (alexandervonhumboldt2):

7x=14

OpenStudy (alexandervonhumboldt2):

x=2

OpenStudy (alexandervonhumboldt2):

x=2 y=0

OpenStudy (anonymous):

@AlexanddervonHumboldt2 Ahh, You got me. Im confused.

OpenStudy (alexandervonhumboldt2):

x=2 y=0

OpenStudy (anonymous):

@AlexandervonHumboldt2 Oh sorry, I dident see that. Thanks man :)

OpenStudy (alexandervonhumboldt2):

ok